Selena wants to invest a portion of her $16000 savings at 8% annual interest, and the balance at a safer 5% annual interest. She hopes to earn $1130 in the next year to pay for a Caribbean Cruise. How much money should she invest at 8%?
Answer by mananth(16946)

Investment I 8.00% per annum ---x
Investment II 5.00% per annum ---y

x + y= 16000 ------------------------1
8.00% x + 5.00% y= = $1,130.00
Multiply by 100
8 x + 5 y= = $113,000.00 --------2

Multiply (1) by -8
we get

-8 x -8 y= = -128000.00

Add this to (2)

0 x -3 y= = -$15,000.00

divide by -3

y = $5,000.00 investment at 5.00%
Balance $11,000.00 investment at 8.00%
CHECK
$11,000.00 --------- 8.00% ------- $880.00
$5,000.00 ------- 5.00% ------- $250.00
Total ------------------- $1,130.00